SPECTROSCOPIC SIGNATURES OF BOND BREAKING INTERNAL ROTATION IN HCP

MARK S CHILD, Physical and Theoretical Chemistry Laboratory, South Parks Rd, Oxford, OX1 3QZ, UK.

Changes in the eigenvalue structure in the vicinty of a saddle-point on the potential energy suface are illustrated by semi-classical and quantum mechanical studies on model potential energy surfaces for HCP. The following points are addressed:

(a) The connection between classical periodic orbits and Fermi resonace polyads, and the breakdown of the polyad model as the bending frequency tunes out of 2:1 resonance with the CP stretch^{a}.

(b) The observation of ’quantum mondromy’ in the underlying spherical pendulum model, and its influence of the values of the spectroscopic vibration-rotation parameters, as the H atom approaches the P end of the molecule^{b}.

(c) A possible formulation of the spectroscopic theory at the saddle point in terms of spherical pendulum eigenstates, and the nature of the relevant matrix elements^{c}.

aM. P. Jacbson and M. S. Child, J. Chem. Phys., 114, 250 (2001)

bM. P. Jacobson and M. S. Child, J. Chem. Phys., 114, 262 (2001).

cM. S. Child, M. P. Jacobson and C. D. Cooper, J. Phys. Chem., 105, 8446 (2001).